OpenStudy (lgbasallote):

hmm i think i've got the idea of circular disk and circular ring methods...can someone show me the formula for cylindrical shell and how to use it? no drawings needed...

OpenStudy (lgbasallote):

lol =))

OpenStudy (inkyvoyd):

I'm extremely helpful, right? @dpaInc @KingGeorge @bmp @myininaya @nbouscal @AccessDenied @apoorvk @FoolForMath

OpenStudy (inkyvoyd):

@myininaya , don't you dare escape my tagarama.

OpenStudy (inkyvoyd):

@lalaly @mimi_x3 @Callisto

OpenStudy (lgbasallote):

lol =)))) though i dont think fool does calculus

OpenStudy (inkyvoyd):

@mimi_x3

OpenStudy (anonymous):

And presto, shorts!

OpenStudy (anonymous):

Sorry, I'd help, but I haven't studied integral volume formulas in aaages. Planning to revisit them soon-ish but not yet.

OpenStudy (lgbasallote):

ahh thanks though :) one less from the tagarama will be fun

OpenStudy (lgbasallote):

though it might have something to do with the fact that @myininaya is the only person that knows calculus and is on right now

OpenStudy (inkyvoyd):

@eigenschmeigen

OpenStudy (farmdawgnation):

@inkyvoyd Knock it off. :P

OpenStudy (lgbasallote):

thanks farm :) i wish they come now =_=

OpenStudy (farmdawgnation):

@inkyvoyd We're going to be taking more stabs at performance issues soon. Sorry for the painfulness. :'(

OpenStudy (accessdenied):

Do you have an example problem? I'm trying to recall what the process was, but it's not coming to me. It's been a while since I ever did these types of problems. lol

OpenStudy (lgbasallote):

well there was this one question that dpaInc said could be done with cylindrical shell as well as washer... Find the volume of the solid obtained by rotating the region bounded by \(y = \frac{x^2}{4}\), \(x = 2\) and \(y = 0\) about the y-axis

OpenStudy (accessdenied):

What answer did you get for it using the other method? I may have this yet! D:

OpenStudy (amistre64):

what is your concept of shells?

OpenStudy (lgbasallote):

my concept of shells? like i know \(\int_a^b \pi r^2 h dx\) like that?

OpenStudy (amistre64):